OpenStudy (xobriley):

4. To which subset(s) of the real numbers does the number -76 belong? a) rational *b) irrational c) integers, whole numbers, rational numbers d) integers, rational numbers

563blackghost (563blackghost):

Natural Numbers: 1, 2, 3, 4 etc. Integers: -3, -4, 0, -6 etc. Rational Numbers: -1.7, 0.75, 7.66.... etc. Irrational Numbers: \(\large{\sqrt{3}}\), \(\large{\pi}\), \(\large{-2\sqrt{5}}\) etc.

OpenStudy (xobriley):

Thank you so intergers

563blackghost (563blackghost):

Correct ^^